OpenStudy (anonymous):

solve e^x + 8e^-x = 6

OpenStudy (anonymous):

Let e^x=a than the equation will be a+8/a=6 /*a a^2+8=6a a^2-6a+8=0 (a-4)(a-2)=0 a=4 or a=2 so e^x=4 implies x=ln4 e^x=2 implies x=ln2